Reports 1-1 of 1 Clear search Modify search
MIF (General)
dan.chen - 17:10 Tuesday 24 September 2024 (31112) Print this report
Xarm cativy scan to make more margin of f1, f2 from carrier.

With KTanaka, Ushiba, Michael,

Summary

  • At first, we adjusted PLLY TEMPY_bias to lock the ALS-Y loop.
  • Then, after the initial alignment for the Xarm, we performed the cavity scan, and found the resonant frequencies between carrier and f1, and f2 were not too close.
  • But, we changed the LO frequency by -5Hz to make more margin. (We confirmed that these 3 don't have overlap.)
  • After this work, we took the alignment for Xarm, Yarm, OMC, and PRMI, and locked PRFPMI.
  • Then DC centering and adjustment of WFS signals were performed.

Details

  • PLLY
    • By whatching K1:ALS-Y_BEAT_ERR_OUT_DQ, change the TEMPY_bias to lock ALS_PLLY.
    • CommissioningDock -> LSC -> ALS OVERVIEW -> Left bottom
    • Procedure
      • Request DOWN to ALS_PLLY
      • Change TEMPY_bias with monitoring K1:ALS-Y_BEAT_ERR_OUT_DQ
      • Request PLL_LOCKED to ALS_PLLY guardian
    • TEMPY_bias:
      • -0.440V -> did not work.
      • -0.545V -> Locked
  • Xarm cativy scan
    • Xarm alignment adjustment was performed by Yokozawa-san
    • Ref: klog-30825
      • Request DOWN to Xarm
      • Change Y offset of ETMX TM SET: -18.1 -> -22.1
      • Requested ALS_LOCKED to XARM
      • Changed the gain of FB of MCL to be 1 as soon as posible.
      • With reference to ENGAGED_ALS_CARM state in "LSC_LOCK" guardian, we performed the following steps
        • Turn ON "FM4" in FILT_MCL
        • Turn ON "IN2" in CMS(Common mode survo) IMC survo (CommissioningDock -> Common Mode Servos -> IMC -> K1:IMC-SERVO_INsEN_SW_ON)
        • Turn OFF "FM2" in FILT_MCL
        • Change "IN2" gain to 24 in CMS_CARM survo. (K1:LSC-CARM_SERVO_IN2GAIN)
        • Turn ON "COMBOOST" in CMS_CARM survo. Turn ON 1 -> Turn ON 2.
        • Turn ON "FM5" in FILT_MCL
      • This time, as we want to see are f1 and f2, we did not need to change the laser power like what Hirose-san's group did.
      • Change the offset of SUMMING SURBO from -27000 to +27000 with 300 seconds RAMP time.
        • Open SUMMING NODE CMS
        • Right click -> Execute -> FB screen -> click SLOWOUT_CALI
        • Set 300s to the RAMP time
        • Change the offset from -27000 to +27000
      • Monitor the flash with K1:LSC-TR_IRX_NORM_OUT_DQ, K1:ALS-SUM_OFS_SLOWOUT_CALI_OUT_DQ, and camera TMSX_IR
        • from 27000 to -27000
        • 1st: TEM00 (offset: 25193)
        • 2nd: TEM01 (offset: 11405)
        • 3rd: TEM00 (offset: 5363)
        • 4th: TEM01 (offset: -8268)
        • 5th: TEM00 (offset: -14155)
        •  
    • Make the alignment back: Change Y offset of ETMX TM SET from -22.1 to -18.0
    • Scaned it again
      • Figures: fig1, fig2
    • We change the modulation frequency.
      • 5.624 369 113 MHz -> 5.624 364 113 MHz
    • Check each peaks
      • df between TEM00 and TEM01 of carrier: 5669.5 [counts]
      • df between TEM00 and TEM01 of carrier: 13779.6
      • df between TEM00 and unknown of RF45: 1700.8
      • df between TEM00 and unknown of RF45: 2466.2
      • df between TEM00 and unknown of RF45: 3394.2
      • df between TEM00 and unknown of RF45: 4159.6
      • We can not see any overlap. So let's use this modulation freq.
      • Figures: fig3, 4
  • Alignemnt adjustment
    • Adjustment of ETMY alignment
      • State at IRY_LOCKED
      • P: -30.2 -> -34.1
      • Y: 3.9 -> 3.1
    • Adjusument of BS
      • State at ALIGNING_YARM
      • P: 10.7 -> 14.7
      • Y: -14.7 -> -16.3
    • OMC
      • Close GRX and Close GRY
      • Requested IRX_LOCKED
      • Requested ALIGNING_TO_OMC
      • Open he GRX, GRY shutters.
    • PRMI
      • Change total(deg) to 176.
      • Requested ALIGNING_PRMI
      • Aalignment adjustment for PRM
        • State: requesting ALIGNING_PRMI
        • See POP_PDA2_RF90_I_ERR_DQ
        • P: 39.0 -> 2.0
        • Y: -56.6 -> -56.6
  • Try lock PRFPMI
    • Requested ENGAGE_WFSDC -> locked.
  • WFS alignment adjustment
    • DC centering
      • Use REFL_WFS picomotors
    • Change set point on ASC_SERVO_VIEW_ARM.adl
      • K1ASC-DHARD_P
      • K1ASC-CHARD_P
      • K1ASC-DHARD_Y
      • K1ASC-CHARD_Y
Images attached to this report
Comments to this report:
takafumi.ushiba - 21:39 Tuesday 24 September 2024 (31113) Print this report

Additional information:

During the summer holidays, it was reported that TEM01 resonance and f1 sideband resonance seem too close (klog30825).
So, our purpose of this measurement is to confirm resonance of carrier TEM01 and f1 sideband have enough separation now.
Since overlap of these two resonances might affect ASC signals, it is important to confirm if these two resonances are overlapped or not.

Figure 1 shows the X arm transmission, REFL RF45, REFL RF17, and offset on SUM CMS.
Left and right cursor show the time when TEM01 and f1 sideband were flashed, respectively.
According to the SUM CMS offset vaue difference (70 cnts) and calibration factor from SUM CMS offset to laser frequency (klog21766: 1.2588 Hz/cnt), frequency difference when resonating TEM01 and f1 sideband is 88 Hz.

Since HWHM of arm cavity is 18 Hz, f1 sideband is about 5*HWHM far from TEM01 resonances.
So, current condition seems not so bad.

Images attached to this comment
Search Help
×

Warning

×